# Position Paper: Quo Vadis, Unsupervised Time Series Anomaly Detection?

### v1

@@ Abstract
The current state of machine learning scholarship in Timeseries Anomaly Detection (TAD) is plagued by the persistent use of flawed evaluation metrics, inconsistent benchmarking practices, and a lack of proper justification for the choices made in novel deep learning-based model designs. Our paper presents a critical analysis of the status quo in TAD, revealing the misleading track of current research and highlighting problematic methods, and evaluation practices. Our position advocates for a shift in focus from pursuing only the novelty in model design to improving benchmarking practices, creating non-trivial datasets, and placing renewed emphasis on studying the utility of model architectures for specific tasks. Our findings demonstrate the need for rigorous evaluation protocols, the creation of simple baselines, and the revelation that state-of-the-art deep anomaly detection models effectively learn linear mappings. These findings suggest the need for more exploration and development of simple and interpretable TAD methods. The increment of model complexity in the state-of-the-art deep-learning based models unfortunately offers very little improvement. We offer insights and suggestions for the field to move forward.
@@ 1 Introduction
Although it is well intended, this line of research has never provided evidence of the necessity of deep-learning, which has been challenged namely in Audibert et al. (2022). The state-of-the-art (SOTA) deep-learning approaches proceeded to introduce models of increased complexity using questionable validation processes. Those processes involve unsuitable benchmark datasets (Wu & Keogh, 2022) and, most harmful to this field, the use of flawed evaluation protocols (Kim et al., 2022). The protocol which introduced the most pitfalls is the point adjustment (PA) applied on the point-wise F1 score which practically favors noisy predictions. It was gradually introduced in a series of papers (Xu et al., 2018; Audibert et al., 2020; Shen et al., 2020; Su et al., 2019c) with the original intention of calibrating the anomaly detection threshold on a hold-out dataset, but it was subsequently demonstrated in Kim et al. (2022) that uniformly random predictions outperform SOTA methods and their performance tends to one as the average length of the anomalies increases. Although using the standard F1 score without point-adjust avoids those pitfalls, it still leaves a gap by only focusing on point-wise time-stamp level detection versus anomaly instance level detection, which led to the introduction of new complementary range-based metrics such as the ones in Tatbul et al. (2018) and Wagner et al. (2023).
The goal of this paper is to guide the TAD community towards more meaningful progress through rigorous benchmarking practices and a focus on studying the utility of their models by drawing useful but simple baselines. We achieve this with the following contributions: 1.) We introduce simple and effective baselines [^1] Code: https://github.com/ssarfraz/QuoVadisTAD and demonstrate that they perform on par or better than the SOTA methods, thus challenging the efficiency and effectiveness of increasing model complexity to solve TAD problems. 2.) We reinforce this position by reducing trained SOTA models to linear models which are distillations of them but still perform on par. Thus from the point of view of the TAD task on the current datasets, those models perform roughly a linear separation of the anomalies from the nominal data.
@@ 2 Related Work
Aside from the anomaly detection approaches, many efforts has been put in creating useful anomaly detection benchmarks. Some recent studies, for instance (Wu & Keogh, 2022) have shown how some of these datasets may suffer from potential flaws, such as triviality, unrealistic density of anomaly, or mislabeling.
@@ 3.1 Preliminaries
The error vectors $\mathbf{E}$ estimated from any of the methods provide a measure of the deviation of the test features from normality. Normalization of error vectors sometimes is necessary before detecting anomalies due to variations in error behavior across sensors. Two normalization methods are often used: scaling using robust statistics such as median and inter-quartile range (Deng & Hooi, 2021) and scaling using mean and standard deviation. The choice of normalization approach can impact anomaly detection accuracy, and careful consideration should be given to the selected method. The impact of error vector normalization on datasets is demonstrated in the appendix through an ablation study in section A.2.1. Once the error vectors are normalized, the final output is a measure of the vector sizes. Given that we are working on the anomaly detection scenario, the most fitting metric is $L^{\infty}$ which computes the largest absolute error between the different sensors, $\|\mathbf{e}_{t}\|_{\infty}=\max\limits_{i\leq F}\{|e_{t}^{i}|\}$.
@@ 3.2 Proposed simple and effective baselines
Figure 1: Proposed simple neural-network baselines
@@ 3.4 Univariate time series representation
Univariate time series data consist of a single observation at each timestamp, and most deep-learning methods designed for multivariate data are not directly applicable. Consequently, the most effective approaches for analyzing univariate data are typically focused on identifying unusual subsequences, or discords, within the time series. State-of-the-art discord discovery methods (Nakamura et al., 2020) focus on optimizing the complexity and parameters of such methods, which typically involve comparing windowed distances between timestamps. In this work, we use a similar yet effective representation for univariate time series data that allows the discovery of anomalies. Specifically, we represent each timestamp as a vector in $\mathbb{R}^{w+1}$, where $w$ denotes the number of preceding time stamps. This representation can be efficiently computed in a sliding window fashion and has linear time complexity, making it efficient for practical use. Furthermore, we demonstrate that the impact of the window size $w$ on performance is relatively low and a small fixed window of $w=4$ suffice for the considered univariate datasets.
@@ 4 Analysis
Secure Water Treatment (SWaT)(Mathur & Tippenhauer, 2016) and Water Distribution (WADI)(Ahmed et al., 2017) datasets: contain sensor measurements of a water treatment test-bed. The WADI dataset demonstrates the inconsistency in reporting performance comparisons in the TAD literature. The complete set of WADI contains 127 sensors (denoted as WADI-127 in our study). However, some recent methods (Tuli et al., 2022; Deng & Hooi, 2021; Kim et al., 2022; Chen et al., 2021) use a specific subset of sensors when making comparisons without specifying the exact used sensors nor the reasons for such selection. Furthermore, in many cases, the selected subsets are inconsistent among competing methods. In order to provide a fair overview of this impact on performance, we conducted our experiments on all the 127 WADI sensors (denoted as WADI-127) and on the subset of 112 sensors used in some recent studies (Deng & Hooi, 2021) (denoted as WADI-112), separately.
@@ 4.1 Model setup
Hyperparameter sensitivity: Most of the simple baselines don’t have tunable hyperparameters. The only exceptions are the projection dimension of the PCA method and the sliding window for univariate series. We have included their ablations in the appendix sections A.2.3 and  A.2.2. We trained our neural network baseline models using the same hyperparameters as stated above on all multivariate datasets. The purpose of this analysis was to demonstrate that even with basic hyperparameters, these simple neural networks can achieve comparable performance to SOTA deep learning models. The fact that the hyperparameters of the SOTA models were optimized for each respective datasets, while the simple NN baseline models used the same set of hyperparameters, highlights less reliance on dataset-specific tuning.
@@ 4.2 Model performance overview
First, it is evident that all methods have higher scores on the predominantly used point-adjusted $F1_{PA}$ metric including the random prediction which performs better in almost all comparisons. This artificial advantage created by point-adjust is not present on the pure F1 score protocols which do not favour noisy random predictions. On both standard point-wise $F1$ and range-wise $F1_{T}$ metrics, the simple baselines such as PCA reconstruction error performs best on all datasets while other baselines such as 1-NN distance and L2-norm are often very close to the best performing methods. Furthermore, the NN-baselines in most cases outperform the more complex SOTA deep models which are build using these as basic building blocks. This is a strong evidence that the complicated solutions introduced to solve the TAD task do not provide a benefit compared to such simple baselines. Finally, one can notice the interplay between the point-wise and range-wise metrics. In datasets like SWAT, where there is a small number of long anomaly intervals, the $F1$ score is much higher than the $F1_{T}$ score, on noisy datasets with more consistent anomaly lengths, like WADI_127, $F1_{T}$ is tendentially higher, while on cleaner datasets with frequent short anomalies, like univariate UCR datasets, the two scores are comparable.
@@ 4.3 Analysis of the deep models learned function
GDN (Deng & Hooi, 2021) 0.810 0.808 0.570 0.543
Table 4: Linear approximation of complex models on two datasets. Orig: original model Line: linear approximated mode. Performance is reported on the standard $F1$ score.
@@ 5 Quo vadis
Aside from exposing the limitations of these methods, we provide a comprehensive set of simple benchmarks which can help re-start investigations in TAD starting on a solid baseline. We think that those methods will pinpoint which anomalies are easy to detect and which ones are the challenging ones that should be detected if any progress is to be made. This is further reinforced by the fact that there seem to be a high agreement between detected and undetected anomalies between all methods investigated. We provide an analysis of this agreement in the appendix section A.1.
Furthermore, we believe that evaluation using both point-wise and range-wise methods will help better compare methods and identify their strengths and weaknesses. We hope our work will help improve the research efforts on TAD by triggering focus on the introduction of new and richer datasets, increasing awareness of limitations of current evaluation protocols, and encouraging caution in the premature adoption of complex tools for the task.
@@ Appendix A Appendix
    Analysis: We analyze the agreement on the detected anomalies between the different models (Figures 3(a) and 3(b)).
@@
    • Impact of normalization: We show the impact of the normalization on the prediction scores (Table 5).
@@
    • Ablation PCA: We show the impact of the PCA dimension used in the PCA reconstruction error method on multivariate datasets (Figure 5).
@@ A.1 Analysis of model agreement on the detected anomalies
In Figures 3(a) and 3(b) the matrices of the agreements between all models and the ground truth are displayed for the SWAT and WADI-112 datasets. In both cases, the agreement between different models is much higher compared to the agreement to the ground truth, indicating that the models learn to recognize similar anomalies. Only the GDN model and even more the PCA Error baseline seem to have a comparably higher agreement with the ground truth.
Those observations lead us to believe that the current datasets used in TAD are, in some sense, simultaneously too hard and too easy. The fact that so many complex deep learning architectures have been developed to tackle the hard anomalies in those datasets, but failed, is unsatisfactory. What is even more worrisome though, is the fact that almost all the recent deep-learning based methods use the point-adjust (PA) post-processing step often without clearly stating this. Under this evaluation these models implicitly optimize for near random predictions where their high performance is used as evidence of their proposed model’s utility. An example of this trend presented at recent leading ML venues is (Xu et al., 2022; Li et al., 2023). Another example is mismatched evaluation metrics i.e., using PA and directly comparing their results to other methods which were scored without it. One such case is GTA (Chen et al., 2021) using PA in evaluation and compares to GDN (Deng & Hooi, 2021) which doesn’t. Similar examples can be found in dataset discrepancy with the introduction of a separate version of the dataset which uses a subset of the sensors (Feng & Tian, 2021; Deng & Hooi, 2021; Chen et al., 2021) and results in higher scores. We hope that the next research steps in TAD will place more focus on the introduction of richer datasets, proper benchmarking, and more rigorous validation of their proposed model design choices.
@@ A.2 Additional evaluations/ablations
A.2.1 Impact of normalization
SWAT WADI-112
None Mean-STD Median-IQR None Mean-STD Median-IQR
F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC
Table 5: Impact of normalization on scores. We show our two best baselines and previously published best-performing GDN and TranAD methods for comparison.
Anomaly detection methods for multivariate datasets often employ normalization and smoothing techniques to address abrupt changes in prediction scores that are not accurately predicted. However, the choice of normalization method before thresholding can impact performance on different datasets. In Table 5, we compared the performance with and without normalization. We consider two normalization methods - mean-standard deviation and median-IQR - on two datasets. Our analysis shows that median-IQR normalization with smoothing, which is also utilized in the GDN (Deng & Hooi, 2021) method, improves performance, especially on noisier datasets such as WADI. In the main paper, we have presented the best performance achieved by each method, including our baselines and considered state-of-the-art models, using either mean-standard deviation or median-IQR normalization, whichever is applicable.
A.2.2 Ablation window size for Univariate data
@@
A.2.3 Ablation PCA: impact of dimension
Figure 5: F1 score as a function of the PCA dimension for the PCA Error method, evaluated on the SWAT and WADI_127 datasets.
A.2.4 NN-baselines: reconstruction vs forecasting mode
@@
A.2.5 Detailed performance comparison
@@
A.3 Performance of our simple baselines on SMAP and MSL datsets

### v2

# Position: Quo Vadis, Unsupervised Time Series Anomaly Detection?

@@ Abstract
The current state of machine learning scholarship in Timeseries Anomaly Detection (TAD) is plagued by the persistent use of flawed evaluation metrics, inconsistent benchmarking practices, and a lack of proper justification for the choices made in novel deep learning-based model designs. Our paper presents a critical analysis of the status quo in TAD, revealing the misleading track of current research and highlighting problematic methods, and evaluation practices. Our position advocates for a shift in focus from solely pursuing novel model designs to improving benchmarking practices, creating non-trivial datasets, and critically evaluating the utility of complex methods against simpler baselines. Our findings demonstrate the need for rigorous evaluation protocols, the creation of simple baselines, and the revelation that state-of-the-art deep anomaly detection models effectively learn linear mappings. These findings suggest the need for more exploration and development of simple and interpretable TAD methods. The increment of model complexity in the state-of-the-art deep-learning based models unfortunately offers very little improvement. We offer insights and suggestions for the field to move forward.
@@ 1 Introduction
Although it is well intended, this line of research has never provided evidence of the necessity of deep-learning, which has been challenged namely in Audibert et al. (2022). The state-of-the-art (SOTA) deep-learning approaches proceeded to introduce models of increased complexity using questionable validation processes. Those processes involve unsuitable benchmark datasets (Wu & Keogh, 2022) and, most harmful to this field, the use of flawed evaluation protocols (Kim et al., 2022). The protocol which introduced the most pitfalls is the point adjustment (PA) applied on the point-wise F1 score which practically favors noisy predictions. It was gradually introduced in a series of papers (Xu et al., 2018; Audibert et al., 2020; Shen et al., 2020; Su et al., 2019c) with the original intention of calibrating the anomaly detection threshold on a hold-out dataset, but it was subsequently demonstrated in Kim et al. (2022) that uniformly random predictions outperform SOTA methods and their performance tends to one as the average length of the anomalies increases. Although using the standard F1 score without point-adjust avoids those pitfalls, it still leaves a gap by only focusing on point-wise time-stamp level detection versus anomaly instance level detection, which led to the introduction of new complementary range-based metrics such as the ones in Tatbul et al. (2018), Wagner et al. (2023).
The goal of this paper is to guide the TAD community towards more meaningful progress through rigorous benchmarking practices and a focus on studying the utility of their models by drawing useful but simple baselines. We achieve this with the following contributions: 1.) We introduce simple and effective baselines and demonstrate that they perform on par or better than the SOTA methods, thus challenging the efficiency and effectiveness of increasing model complexity to solve TAD problems. 2.) We reinforce this position by reducing trained SOTA models to linear models which are distillations of them but still perform on par. Thus from the point of view of the TAD task on the current datasets, those models perform roughly a linear separation of the anomalies from the nominal data.
Our code [^1] Code: https://github.com/ssarfraz/QuoVadisTAD is available on GitHub to easily run the baselines and benchmarks.
@@ 2 Related Work
Aside from the anomaly detection approaches, many efforts has been put in creating useful anomaly detection benchmarks. Some recent studies, for instance (Wu & Keogh, 2022) have shown how some of these datasets suffer from potential flaws, such as triviality, unrealistic density of anomaly, or mislabeling.
@@ 3.1 Preliminaries
The error vectors $\mathbf{E}$ estimated from any of the methods provide a measure of the deviation of the test features from normality. Normalization of error vectors sometimes is necessary before detecting anomalies due to variations in error behavior across sensors. Two normalization methods are often used: scaling using robust statistics such as median and inter-quartile range (Deng & Hooi, 2021) and scaling using mean and standard deviation. The choice of normalization approach can impact anomaly detection accuracy, and careful consideration should be given to the selected method. The impact of error vector normalization on datasets is demonstrated through an ablation study in section 4.4. Once the error vectors are normalized, the final output is a measure of the vector sizes. Given that we are working on the anomaly detection scenario, the most fitting metric is $L^{\infty}$ which computes the largest absolute error between the different sensors, $\|\mathbf{e}_{t}\|_{\infty}=\max\limits_{i\leq F}\{|e_{t}^{i}|\}$.
@@ 3.2 Proposed simple and effective baselines
Figure 1: Proposed simple neural-network baselines
@@ 3.4 Univariate time series representation
Univariate time series data consist of a single observation at each timestamp, and most deep-learning methods designed for multivariate data are not directly applicable. Consequently, the most effective approaches for analyzing univariate data are typically focused on identifying unusual subsequences, or discords, within the time series. State-of-the-art discord discovery methods, for instance (Nakamura et al., 2020), focus on optimizing the complexity and parameters of such methods that typically involve comparing windowed distances between timestamps. In this work, we use a similar yet effective representation for univariate time series data that allows the discovery of anomalies. Specifically, we represent each timestamp as a vector in $\mathbb{R}^{w+1}$, where $w$ denotes the number of preceding time stamps. This representation can be efficiently computed in a sliding window fashion and has linear time complexity, making it efficient for practical use. Furthermore, we demonstrate that the impact of the window size $w$ on performance is relatively low and a small fixed window of $w=4$ suffice for the considered univariate datasets.
@@ 4 Analysis
Secure Water Treatment (SWaT) (Mathur & Tippenhauer, 2016) and Water Distribution (WADI) (Ahmed et al., 2017) datasets: contain sensor measurements of a water treatment test-bed. Although SWaT is commonly used as a benchmark in recent publications, it should be noted that its use as a benchmark should be discontinued as it is flawed and unreliable Eamonn Keogh (personal communication, 7 May, 2024), see also (Wagner et al., 2023). The WADI dataset demonstrates the inconsistency in reporting performance comparisons in the TAD literature. The complete set of WADI contains 127 sensors (denoted as WADI-127 in our study). However, some recent methods (Tuli et al., 2022; Deng & Hooi, 2021; Kim et al., 2022; Chen et al., 2021; Feng & Tian, 2021) use a specific subset of sensors when making comparisons without specifying the exact used sensors nor the reasons for such selection. Furthermore, in many cases, the selected subsets are inconsistent among competing methods. In order to provide a fair overview of this impact on performance, we conducted our experiments on all the 127 WADI sensors (denoted as WADI-127) and on the subset of 112 sensors used in some recent studies (Deng & Hooi, 2021) (denoted as WADI-112), separately.
@@ 4.1 Model setup
Hyperparameter sensitivity: Most of the simple baselines don’t have tunable hyperparameters. The only exceptions are the projection dimension of the PCA method and the sliding window for univariate series. We have included their ablations in sections 4.5 and  A.2.1. We trained our neural network baseline models using the same hyperparameters as stated above on all multivariate datasets. The purpose of this analysis was to demonstrate that even with basic hyperparameters, these simple neural networks can achieve comparable performance to SOTA deep learning models. The fact that the hyperparameters of the SOTA models were optimized for each respective datasets, while the simple NN baseline models used the same set of hyperparameters, highlights less reliance on dataset-specific tuning.
@@ 4.2 Model performance overview
First, it is evident that all methods have higher scores on the predominantly used point-adjusted $F1_{PA}$ metric including the random prediction which performs better in almost all comparisons. This artificial advantage created by point-adjust is not present on the pure F1 score protocols which do not favour noisy random predictions. On both standard point-wise $F1$ and range-wise $F1_{T}$ metrics, the simple baselines such as PCA reconstruction error performs better on all datasets while other baselines such as 1-NN distance and L2-norm are often very close to the best performing methods. Furthermore, the NN-baselines in most cases outperform the more complex SOTA deep models which are build using these as basic building blocks. This is a strong evidence that the complicated solutions introduced to solve the TAD task do not provide a benefit compared to such simple baselines. Finally, one can notice the interplay between the point-wise and range-wise metrics. In datasets like SWAT, where there is a small number of long anomaly intervals, the $F1$ score is much higher than the $F1_{T}$ score, on noisy datasets with more consistent anomaly lengths, like WADI_127, $F1_{T}$ is tendentially higher, while on cleaner datasets with frequent short anomalies, like univariate UCR datasets, the two scores are comparable.
@@ 4.3 Analysis of the deep models learned function
GDN (Deng & Hooi, 2021) 0.810 0.808 0.571 0.543
Table 4: Linear approximation of complex models on two datasets. Orig: original model Line: linear approximated mode. Performance is reported on the standard point-wise $F1$ score.
SWAT WADI-112
None Mean-STD Median-IQR None Mean-STD Median-IQR
F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC
Table 5: Impact of normalization on scores. Normalisation of prediction scores before thresholding impacts performance. Performance is reported on the point-wise $F1$ and $AUPRC$ score.
4.4 Ablation: Impact of normalization
Anomaly detection methods for multivariate datasets often employ normalization and smoothing techniques to address abrupt changes in prediction scores that are not accurately predicted. However, the choice of normalization method before thresholding can impact performance on different datasets. In Table 5, we compared the performance with and without normalization. We consider two normalization methods, mean-standard deviation and median-IQR, on two datasets. Our analysis shows that median-IQR normalization, which is also utilized in the GDN (Deng & Hooi, 2021) method, improves performance on noisier datasets such as WADI. In Table 2, we have presented the best performance achieved by each method, including our baselines and considered state-of-the-art models, using either none or one of these normalisation, whichever is applicable.
4.5 Ablation: PCA Error projection dimension
Figure 3: Point-wise F1 score as a function of the PCA dimension for the PCA Error method, evaluated on the SWAT and WADI_127 datasets.
@@ 5 Quo vadis
We cannot stress enough the fact that almost all the recent deep-learning based methods use the point-adjust post-processing step often without clearly stating this. Under this evaluation these models implicitly optimize for near random predictions where their high performance is used as evidence of their proposed model’s utility. An example of this trend presented at recent leading Machine Learning venues is (Xu et al., 2022; Li et al., 2023). Another common malpractice, is the use of mismatched evaluation metrics in tables i.e., applying point-adjust and directly comparing their results to other methods which were scored without it. Similar issues are observed in dataset discrepancies like the introduction of new versions of a dataset which use a subset of the sensors and result in higher scores.
Aside from exposing the limitations of these methods, we provide a comprehensive set of simple benchmarks which can help re-start investigations in TAD starting on a solid baseline. We think that those methods will pinpoint which anomalies are easy to detect and which ones are the challenging ones that should be detected if any progress is to be made. This is further reinforced by the fact that there seem to be a high agreement between detected and undetected anomalies between all methods investigated. We provide an analysis of this agreement in the appendix section A.1. This agreement leads us to believe that the current datasets used in TAD are, in some sense, simultaneously too hard and too easy. The fact that so many complex deep learning architectures have been developed to tackle the hard anomalies in those datasets, but failed, is unsatisfactory, but maybe not unexpected. More comprehensive datasets with a spread spectrum of difficulty in anomalies could provide an incremental improvement path and means of properly comparing methods.
Furthermore, we believe that evaluation using both point-wise and range-wise methods will help better compare methods and identify their strengths and weaknesses.
We hope our work will help improve the research efforts on TAD by triggering focus on the introduction of new and richer datasets, increasing awareness of limitations of current evaluation protocols, and encouraging caution in the premature adoption of complex tools for the task.
@@ Appendix A Appendix
    Analysis: We analyze the agreement on the detected anomalies between the different models (Figures 4(a) and 4(b)).
@@ A.1 Analysis of model agreement on the detected anomalies
In Figures 4(a) and 4(b) the matrices of the agreements between all models and the ground truth are displayed for the SWAT and WADI-112 datasets. In both cases, the agreement between different models is much higher compared to the agreement to the ground truth, indicating that the models learn to recognize similar anomalies. Only the GDN model and even more the PCA Error baseline seem to have a comparably higher agreement with the ground truth.
@@ A.2 Additional evaluations/ablations
A.2.1 Ablation window size for Univariate data
@@
A.2.2 NN-baselines: reconstruction vs forecasting mode
@@
A.2.3 Detailed performance comparison
@@
A.3 Performance of our simple baselines on SMAP and MSL datasets